jonathangh:

Each month, Stacey places a portion of her earnings in a retirement account. Her savings are invested in two companies. For every $9 invested in company A, $2 is invested in company B. Stacey quizzes her children, Krista and Kevin, about how much money she will invest in company A if she invests $24 in company B.

sealake101:

independent variable,be as X, $9 dependent variable, be as S $2 multiplier of x to get the product Y (X = 9 and S = 2) 9 = (y) (2) 9 รท 2 = y 4.5 = y ( S = 24) Change S, while Y remains constant X = (4.5) (24) ok so whats 4.5 x 24?

sealake101:

x=_____ ( that would be your answer)